ID |
Date |
Author |
Type |
Category |
Subject |
7618
|
Thu Oct 25 06:49:49 2012 |
Koji | Update | | my assesment of the folding mirror (passive tip-tilt) situation |
Quote: |
My big question right now is: is the plan to install new dichroics in PR2 and SR2 as well, or just in PR3 and SR3, where the green beams are extracted? I think the answer is no, we only want to install new dichroics in {P,S}R3.
|
Why not? The new dichroic mirrors have more transmission of 1064nm than G&H. Thus it will give us more POP beam and will help locking. |
7617
|
Thu Oct 25 02:10:22 2012 |
Koji | Update | | my assesment of the folding mirror (passive tip-tilt) situation |
The thinner wire has a history that it did not improve the hysteresis (ask Jenne). Nevertheless, it's worth to try.
If you flip the clamp upside-down, you can lift the clamping point up. This will make the gravity restoring torque stronger.
(i.e. Equivalent effect to increasing the mass)
Luckily (or unluckily) the clamp has no defined location for the wire as we have no wire fixture.
Therefore the clamp will grab the wire firmly even without milling. |
7616
|
Thu Oct 25 02:01:15 2012 |
Koji | Update | Computer Scripts / Programs | Phase map summary of LaserOptik mirrors |
Previous phasemap data and analysis for the new 40m COC are summarized on the following page
https://nodus.ligo.caltech.edu:30889/40m_phasemap/
(Use traditional LVC authentication (not albert.einstein))
The actual instance of the files can also be found on nodus below the following directory:
/cvs/cds/caltech/users/public_html/40m_phasemap
The programs for the analysis are found in
/cvs/cds/caltech/users/public_html/40m_phasemap/40m_PRM/mat
The main program is RunThis.m
Basically this program takes ascii files converted from opd by Vision32.
(i.e. You need to go to Downs)
Then the matlab program takes care of the plots and curvature analyses. |
7615
|
Wed Oct 24 22:48:46 2012 |
janosch | Update | Computer Scripts / Programs | Phase map summary of LaserOptik mirrors |
Quote: |
After a long search, I've found a way to finally read and analyze(?) the Wyko opd format data using Image SXM, an image analysis software working only on mac osx.
I am attaching the images (in tiff) and profile plot of all the 6 mirrors.
|
Great, however, unless you can save the images in FITS format, we still need another reader for the opd images. |
7614
|
Wed Oct 24 22:20:24 2012 |
Den | Update | Adaptive Filtering | microphone noise |
Quote: |
We have to change the sample rate and AA filter for the mic channels before going too far with the circuit design.
|
PEM model is running at 64K now. It turned out to be tricky to increase the rate:
- BLRMS are computationally expensive and original pem model did not start at any frequency higher then 16k ( at 16k cpu meter readings were 59/60 ). Also when we go higher then 16k, front-end gives the model less resources. I guess it is assumed that this model is iop and won't need too much time. So in the end I had to delete BLRMS blocks for all channels except for GUR2Z and MIC1.
- Foton files are modified during model compilation: lines with sampling rate and declaration of filters in the beginning of the file are changed only. Sos-representation and commands are the same. I hoped that filter commands will let me change sos-representation quickly. I've opened Foton and saved the file. However, Foton modified commands in such a way that the ratio of poles and zeros to sampling rate is preserved. I guess all filters have to be replaced or this process should be done in another way.
- BLRMS block uses low-pass filters below 0.01 Hz, increasing the sampling rate by a factor of 32 might make calculations incorrect. I'll check it.
We should also increase cut off frequency of the low-pass filter in the microphone pre-amplifier from 2 kHz up to ~20-30 kHz. |
Attachment 1: mic_64k.pdf
|
|
7613
|
Wed Oct 24 20:09:41 2012 |
jamie | Update | | installing the new dirchoic mirros in PR3/SR3 |
When installing the dichroics we need to pay attention to the wedge angle. I didn't, so the ghost beam is currently point up and to the right (when facing the optic). We should think carefully about where we want the ghost beams to go.
I also was using TT SN003, which I believe was being used for PR2. However, I don't think we want to install dichroics in the PR2, and we might want to put all the tip-tilts back in the same spots they were in before. We therefore may want to put the old optic back in SN003, and put the dichroics in SN005 (PR3) and SN001 (SR3) (see 7601). |
7612
|
Wed Oct 24 19:55:06 2012 |
jamie | Update | | my assesment of the folding mirror (passive tip-tilt) situation |
We removed all the folding mirrors ({P,S}R{2,3}) from the IFO and took them into the bake lab clean room. The idea was that at the very least we would install the new dichroic mirrors, and then maybe replace the suspension wires with thinner ones.
I went in to spend some quality time with one of the tip-tilts. I got the oplev setup working to characterize the pointing.
I grabbed tip-tilt SN003, which was at PR2. When I set it up it was already pointing down by a couple cm over about a meter, which is worse than what we were seeing when it was installed. I assume it got jostled during transport to the clean room?
I removed the optic that was in there and tried installing one of the dichroics. It was essentially not possible to remove the optic without bending the wires by quite a bit (~45 degrees). I decided to remove the whole suspension system (top clamps and mirror assembly) so that I could lay it flat on the table to swap the optic.
I was able to put in the dichroic without much trouble and get the suspension assembly back on to the frame. I adjusted the clamp at the mirror mount to get it hanging back vertical again. I was able to get it more-or-less vertical without too much trouble.
I poked at the mirror mount a bit to see how I could affect the hysteresis. The answer is quite a bit, and stochastically. Some times I would man-handle it and it wouldn't move at all. Sometimes I would poke it just a bit and it would move by something like a radian.
A couple of other things I noted:
- The eddy current damping blocks are not at all suspended. The wires are way too think, so they're basically flexures. They were all pretty cocked, so I repositioned them by just pushing on them so they were all aligned and centered on the mirror mount magnets.
- The mirror mounts are very clearly purposely made to be light. All mass that could be milled out has been. This is very confusing to me, since this is basically the entire problem. Why were they designed to be so light? What problem was that supposed to solve?
I also investigated the weights that Steve baked. These won't work at all. The gap between the bottom of the mirror mount and the base is too small. Even the smalled "weights" would hit the base. So that whole solution is a no-go.
What else can we do?
At this point not much. We're not going to be able to install more masses without re-engineering things, which is going to take too much time. We could install thinner wires. The wires that are being used now are all 0.0036", and we could install 0.0017" wires. The problem is that we would have to mill down the clamps in order to reuse them, which would be time consuming.
The plan
So at this point I say we just install the dichroics, get them nicely suspended, and then VERY CAREFULLY reinstall them. We have to be careful we don't jostle them too much when we transport them back to the IFO. They look like they were too jostled when they were transported to the clean room.
My big question right now is: is the plan to install new dichroics in PR2 and SR2 as well, or just in PR3 and SR3, where the green beams are extracted? I think the answer is no, we only want to install new dichroics in {P,S}R3.
The future
If we're going to stick with these passive tip-tilts, I think we need to consider machining completely new mirror mounts, that are not designed to be so light. I think that's basically the only way we're going to solve the hysteresis problem.
I also note that the new active tip-tilts that we're going to use for the IO steering mirrors are going to have all the same problems. The frame is taller, so the suspensions are longer, but everything else, including the mirror mounts are exactly the same. I can't see that they're not going to suffer the same issues. Luckily we'll be able to point them so I guess we won't notice. |
7611
|
Wed Oct 24 18:42:39 2012 |
Manasa | Update | Computer Scripts / Programs | Phase map summary of LaserOptik mirrors |
Quote: |
Raji took the optics over. They were all measured at 0 deg incidence angle, although we will use them at the angles required for the recycling folding mirrors. Here's the summary from GariLynn:
In general all six pieces have a radius of curvature of around -700 meters.
They all fall off rapidly past 40 mm diameter. Within the 40 mm diameter the rms is ~10 nm for most. I can get finer analysis if you have something specific that you want to know.
All data are saved in Wyko format at the following location:
Gari
|
After a long search, I've found a way to finally read and analyze(?) the Wyko opd format data using Image SXM, an image analysis software working only on mac osx.
I am attaching the images (in tiff) and profile plot of all the 6 mirrors. |
Attachment 1: sn1Laseroptik_profile
|
Attachment 2: sn2Laseroptik_profile
|
Attachment 3: sn3Laseroptik_profile
|
Attachment 4: sn4Laseroptik_profile
|
Attachment 5: sn5Laseroptik_profile
|
Attachment 6: sn6Laseroptik_profile
|
Attachment 7: sn1.png
|
|
Attachment 8: sn2.png
|
|
Attachment 9: sn3.png
|
|
Attachment 10: sn4.png
|
|
Attachment 11: sn5.png
|
|
Attachment 12: sn6.png
|
|
7610
|
Wed Oct 24 17:02:01 2012 |
Jenne | Update | Adaptive Filtering | microphone noise |
Quote: |
We have to change the sample rate and AA filter for the mic channels before going too far with the circuit design.
|
To save the mic channels at higher than 2k (which we should do), we either have to move them to a different model, change the rate of the PEM model, or see if you can save data faster than the model runs (which I can't imagine is possible). |
7609
|
Wed Oct 24 15:29:52 2012 |
rana | Update | Adaptive Filtering | microphone noise |
We have to change the sample rate and AA filter for the mic channels before going too far with the circuit design. |
7608
|
Wed Oct 24 14:19:01 2012 |
Jenne | Update | Alignment | Phase map summary of LaserOptik mirrors |
Quote: |
Jamie has arranged for phase map measurements this afternoon, so I will take the 6 dichroic LaserOptik optics over to Downs at 1:15 this afternoon.
Team Jamie+Nic will lead the effort to clamp down dog clamps as placement markers for all 4 in-vac passive TTs, and then pull all 4 TTs out of the chambers. They plus Den will move the TTs to the Cleanroom, and will start to install the new pitch alignment hardware.
When I return with the optics, we will install them in the TTs and re-balance them. Then we can put them back in the chambers and get back to work on alignment.
After we re-install the TTs, we will need to check the leveling of all 3 corner tables, just to be sure.
|
Raji took the optics over. They were all measured at 0 deg incidence angle, although we will use them at the angles required for the recycling folding mirrors. Here's the summary from GariLynn:
In general all six pieces have a radius of curvature of around -700 meters.
They all fall off rapidly past 40 mm diameter. Within the 40 mm diameter the rms is ~10 nm for most. I can get finer analysis if you have something specific that you want to know.
All data are saved in Wyko format at the following location:
Gari
|
7607
|
Wed Oct 24 14:15:34 2012 |
Ayaka | Update | Adaptive Filtering | microphone noise |
Previous results
I am measuring the noise level of the microphones. The circuit does not seems to limit their sensitivities but the circuit's noise seems to be different from other channels.
Measurement
I measured the circuit noise of all 6 channels. (input open)
(mic_open.png)
The noise level is about 10 times different from the others.
Comparing the acoustic signal, microphone+circuit noise, and ADC noise;
(mic_noise.png)
- blue; acoustic signal
- green; microphone+circuit noise
- red; circuit (the data was not took simultaneously.)
- sky blue; ADC noise
To do
I will remake the circuit though the circuit does not limit the sensitivity. I would like to make sure that the circuit does not affect badly and to make the circuit noise level the same.
At the same time, I will get the PMC control signal and see coherence between it and acoustic sound. |
Attachment 1: mic_open.png
|
|
Attachment 2: mic_noise.png
|
|
7606
|
Wed Oct 24 11:49:07 2012 |
Jenne | Update | Alignment | Game plan for the day |
Jamie has arranged for phase map measurements this afternoon, so I will take the 6 dichroic LaserOptik optics over to Downs at 1:15 this afternoon.
Team Jamie+Nic will lead the effort to clamp down dog clamps as placement markers for all 4 in-vac passive TTs, and then pull all 4 TTs out of the chambers. They plus Den will move the TTs to the Cleanroom, and will start to install the new pitch alignment hardware.
When I return with the optics, we will install them in the TTs and re-balance them. Then we can put them back in the chambers and get back to work on alignment.
After we re-install the TTs, we will need to check the leveling of all 3 corner tables, just to be sure. |
7605
|
Wed Oct 24 09:15:12 2012 |
Steve | Update | VAC | vacuum rack 24V ps was current limited |
Quote: |
Quote: |
Quote: |
Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr. All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open. One of the roughing pumps was also turned on. Other stuff we didn't notice? Bad.
|
Several of the suspensions were kicked pretty hard (600+ mV on some sensors) as a result of this quick vent wind. All of the suspensions are damped now, so it doesn't look like we suffered any damage to suspensions.
|
CLOSE CALL on the vacuum system:
Jamie and I disabled V1, VM2 and VM3 gate valves by disconnecting their 120V solenoid actuator before the swap of the VME crate.
The vacuum controller unexpectedly lost control over the swap as Jamie described it. We were lucky not to do any damage! The ion pumps were cold and clean. We have not used them for years so their outgassing possibly accumulated to reach ~10-50 Torr
I disconnected_ immobilized and labelled the following 6 valves: the 4 large ion pump gate valves and VC1, VC2 of the cryo pump. Note: the valves on the cryo pump stayed closed. It is crucial that a warm cry pump is kept closed!
This will not allow the same thing to happen again and protect the IFO from warm cryo contamination.
The down side of this that the computer can not identify vacuum states any longer.
This vacuum system badly needs an upgrade. I will make a list.
|
While I was doing the oil change of the roughing pumps I accidentally touched the 24 V adjustment knob on the power supply.
All valve closed to default condition. I realized that the current indicator was red at 0.2A and the voltage fluctuated from 3-13V
Increased current limiter to 0.4A and set voltage to 24V I think this was the reason for the caos of valve switching during the VME swap.
|
Attachment 1: currentlimited24V.jpg
|
|
7604
|
Wed Oct 24 01:02:10 2012 |
Koji | Update | SUS | PITCH damping needed |
Wow... This is even more complicated than the original "Y" design... |
7603
|
Tue Oct 23 18:21:21 2012 |
Jenne | Update | Alignment | Power supply at OMC removed |
Quote: |
Quote: |
Manasa and Raji hooked up HV power supplies to the PZTs and set them to the middle of their ranges (75 V).
|
[Raji, Manasa]
The high-voltage power supply from the OMC was removed to replace one of the PZT power supplies. The power supply terminals were connected to the rear connection ports as per instructions from the manual (TB1 panel: port 3 - (-)OUT and port7 - (+)OUT). They were both switched on and set to deliver (75V) to the PZTs.
|
This means that the low voltage dual supply which was wired in series (so could supply a max of 63V = 2*31.5V) has been replaced with the OMC power supply. This is okay since we haven't turned on the OMC PZTs in a long, long time. This is *not* the power supply for the output pointing PZTs. When she says "both", she means the new HV supply, as well as the HV supply that was already there, so both pitch and yaw for PZT2 are being supplied with 75V now. |
7602
|
Tue Oct 23 18:18:29 2012 |
Jenne | Update | SUS | PITCH damping needed |
Quote: |
Quote: |
Too bad - I thought it would at least give a little damping. Since we want the viscous-like energy loss to be ~49x larger, we need to have the field modulation in the damper (not dumper) increase by ~7.
|
I've made SolidWorks models of damping bracket and eddy current disk. They will me manufactured and used instead of old ones. New bracket will be mounted in exactly the same place where the old one was. Drawings might not be complete but all dimensions are in the models so we can fix drawing tomorrow before going to machine shop.
I think we can use ring magnets for passive damping. Then we won't have the vent problem. I've found some at K&J Magnetics, we can get them any time. Magnets are Ni-Cu-Ni (fine for vacuum?) Diameter is 3/8'' with advertised tolerence 0.004'', so they should fit the holes.
|
Den mentioned that the disks will have threaded holes, and that he has made a note to that effect on the paper copy of the drawing that he will bring to Mike at the shop. Also, all threaded holes in the new plate are marked on the paper copy. |
7601
|
Tue Oct 23 18:12:18 2012 |
Jenne | Update | Alignment | Tip tilt wires - the truth |
Quote: |
At this point I cried foul. This is not an acceptable situation. Very little stimulation to the tip-tilts can repoint the beam inside the PR cavity.
Steve says that the TT weights, which will attach to the base of the TT mirror mounts and should help keep the mirrors vertical and not hysteretic, are being baked now and should be available tomorrow.
|
We also wrote down the serial numbers (top center of each TT, inscribed by hand) for what tip tilt is installed where. I then went through the elog to determine which TT was suspended with what kind of wire (thick or thin). Summary: all installed tip tilts have thick wire, 0.0036" diameter.
As noted in elog 3295, we had found that there was similar hysteresis whether we used the thick or the thin wire, so we had decided not to go back and re-suspend every optic.
Also, since we will redo the pitch balance tomorrow with the new hardware tomorrow, I think we should put in the new LaserOptik mirrors at the same time. We have not yet gotten phase maps of them, but we might as well do this rebalancing once, rather than twice.
As-installed tip tilt list
Serial number |
Installed as |
Wire thickness |
Notes, elog reference |
001 |
SR 3 |
0.0036" |
See elog 3437 |
002 |
SR 2 |
0.0036" |
See elog 3295 |
003 |
PR 2 |
0.0036" |
No elog, but inferred since there were 4 with thick wire, and #004 is the thin wire one. Elog 3437 has notes on the 4 thick, 1 thin situation. |
004 |
spare, dirty |
originally 0.0017", but looks redone with thicker wire |
See elog 3295 |
005 |
PR 3 |
0.0036" |
Was supposed to be spare according to elog 3437, but was installed. See elog 3437 |
|
7600
|
Tue Oct 23 17:41:20 2012 |
Manasa | Update | Alignment | Power supply at OMC removed |
Quote: |
Manasa and Raji hooked up HV power supplies to the PZTs and set them to the middle of their ranges (75 V).
|
[Raji, Manasa]
The high-voltage power supply from the OMC was removed to replace one of the PZT power supplies. The power supply terminals were connected to the rear connection ports as per instructions from the manual (TB1 panel: port 3 - (-)OUT and port7 - (+)OUT). They were both switched on and set to deliver (75V) to the PZTs.
|
7599
|
Tue Oct 23 17:30:33 2012 |
jamie, nic, jenne, raji, manasa | Update | Alignment | Initial attempts to fix IFO alignment |
We went into the vertex today to see about fixing the alignment. The in-air access connector is in place, and we took heavy doors off of BS, ITMY, and ETMY chambers.
We started by looking at the pointing from the PZTs. Manasa and Raji hooked up HV power supplies to the PZTs and set them to the middle of their ranges (75 V).
We installed a target on the BS cage, and new "free standing" targets made special by Steve for the SOSs on ITMY and ETMY.
Using a free-standing aperture target we looked at the beam height before PZT2. It was a little high, so we adjusted it with PZT1. Once that was done we looked at the beam height at PR2, and adjusted that height with PZT1.
We then tried to use the hysteresis in PR2 to adjust the beam height at ITMY. Pushing just a little bit at the top or bottom of PR2 would repoint the beam in pitch. This sort of works, but it's stupid. Using this method we got the beam more or less centered vertically at ITMY.
We moved on to ETMY with the idea that we would again use the hysteresis in PR3 to get the vertical pointing to the ETM correct. This was a good demonstration of just how stupid the tip-tilts really are. Just touching slightly at the top or bottom or PR3 we could completely change the pointing at ETMY, by mili-radians (~4 cm over 40m).
At this point I cried foul. This is not an acceptable situation. Very little stimulation to the tip-tilts can repoint the beam inside the PR cavity.
Steve says that the TT weights, which will attach to the base of the TT mirror mounts and should help keep the mirrors vertical and not hysteretic, are being baked now and should be available tomorrow. We therefore decided to stop what we were doing today, since we'll have to just redo it all again tomorrow once the weights are installed.
|
7598
|
Tue Oct 23 17:12:30 2012 |
Den | Update | SUS | PITCH damping needed |
Koji and Steve pointed out that previous design of a damping bracket was a bit complicated to manufacture. So I made it simpler and also added a tap hole for original yaw damping. We'll give drawing to Mike in the machine shop tomorrow morning.
I've purchased K&J magnets for eddy current damping, they should be here in 2 days. |
Attachment 1: simple_drawing.PDF
|
|
7597
|
Tue Oct 23 16:18:03 2012 |
Steve | Update | VAC | Roughing pumps on for oil change |
Leybold D30AQuote: |
PR1, PR2 and RP3 turned on for warming up for oil change. Oil changed with 3.2L of MVT-19 fluid in each. This substitute for HE-175 will be used next time.
To finish this job tomorrow: 1, check oil creeping upstream 2, change air filter of air purge if pressure drops <350 mTorr 3, measure venting time of pump
|
. Leybold D30A manual is here. Exhaust filter traps were drained. No oil creeping was found. The pump venting time is < 1 minute.
PR-2 needs new secu-valve. PR1 & 2 are in excellent condition. |
Attachment 1: oilchange.jpg
|
|
7596
|
Tue Oct 23 10:24:42 2012 |
Ayaka | Update | Adaptive Filtering | microphone noise |
Quote: |
Quote: |
I will do some experiments on acoustic noise canceling during my stay.
Now I am planning to cancel acoustic noise from PMC and see how the acoustic noise work and how we should place microphones.a
First, I measured the noise in microphones and its circuit.

-blue, green, red, solid lines; microphone signals
-blue, green, red, dashed lines; un-coherent noise in signals
-yellow, black, solid lines; circuit noise (signal input is open, not connected to the microphones)
We can see the acoustic signal above 1 Hz, and the circuit does not seem to limit its sensitivity. But I do not know why yellow and black is so different. I will check it tomorrow.
|
Hi, Ayaka. It would be good if you could give a little bit more detail about this plot:
- What exactly are the "signals"? Are you making a sound somehow? If so, what is producing the sound? What is it's spectrum?
- Are the blue/green/red traces from three different microphones?
- Coherence usually implies a comparison between two signals. Is something being compared in the dashed traces?
- Are the yellow and black traces from different amplifiers?
- What are the units of the Y axis?
|
Sorry for my poor explanation.
I measured this by the same way as you measured the instrumental noise of seismometers.
I put the three microphones at the same place so that the three can hear the same sound. I did not make any sounds, just put them in the lab.
The signals from microphones are all amplified by the circuit.
And I took the correlations of each signals and two others and got the noise (dashed lines) by subtracting the correlated signal from the original signal.
So,
-The signal is the acoustic sound in the lab, amplified by the circuit.
-Three lines are from three different microphones.
-Dashed lines are subtraction of coherent signal from the original.
-Yellow and black lines are from different amplifiers in the same circuit box. The circuit has 6 channels.
-I did not calibrate the signals I got by DTT since I do not know the calibration factor now. It is just the number I got from the real time system.
|
7595
|
Tue Oct 23 09:37:33 2012 |
Steve | Update | SUS | passive TT adjustment correction |
Quote: |
Tip Tilt pitch adjustment on existing-in vacuum suspension. This can be added by a simple installation of a 1.25" long 2-56 threaded rod with nuts.
|
We should check that their sus wire diameter are 0.0017" All 2-56 hardware are in and Bob is cleaning them. |
Attachment 1: IMG_1745.JPG
|
|
7594
|
Tue Oct 23 08:45:00 2012 |
Steve | Update | VAC | vent day 3 |
Quote: |
Quote: |
Quote: |
PSL shutter closed, manual block in place, HV turned off. P1 is at 200 Torr now. Jenne is taking over here.
|
Valves closed, 500 torr. Steve will finish off Monday morning, then we'll take off doors and get to work.
|
We are almost at atm. P1 750 Torr, We are slowly reaching equilibrium.
|
The access connector was removed, BS and ITMY chambers were opened yesterday.
New one piece aluminum shell access connector installed. Jenne lowered the PSL out put into MC to ~100mW |
Attachment 1: ventd4.png
|
|
Attachment 2: IMG_1743.JPG
|
|
7593
|
Tue Oct 23 01:46:53 2012 |
Den | Update | SUS | PITCH damping needed |
Quote: |
Too bad - I thought it would at least give a little damping. Since we want the viscous-like energy loss to be ~49x larger, we need to have the field modulation in the damper (not dumper) increase by ~7.
|
I've made SolidWorks models of damping bracket and eddy current disk. They will me manufactured and used instead of old ones. New bracket will be mounted in exactly the same place where the old one was. Drawings might not be complete but all dimensions are in the models so we can fix drawing tomorrow before going to machine shop.
I think we can use ring magnets for passive damping. Then we won't have the vent problem. I've found some at K&J Magnetics, we can get them any time. Magnets are Ni-Cu-Ni (fine for vacuum?) Diameter is 3/8'' with advertised tolerence 0.004'', so they should fit the holes. |
Attachment 1: Mirror_Holder_ECD_NEW_DRAWING.PDF
|
|
Attachment 2: TT_DAMPING_DRAWING.PDF
|
|
7592
|
Tue Oct 23 00:51:41 2012 |
Jamie | Update | Adaptive Filtering | microphone noise |
Quote: |
I will do some experiments on acoustic noise canceling during my stay.
Now I am planning to cancel acoustic noise from PMC and see how the acoustic noise work and how we should place microphones.a
First, I measured the noise in microphones and its circuit.

-blue, green, red, solid lines; microphone signals
-blue, green, red, dashed lines; un-coherent noise in signals
-yellow, black, solid lines; circuit noise (signal input is open, not connected to the microphones)
We can see the acoustic signal above 1 Hz, and the circuit does not seem to limit its sensitivity. But I do not know why yellow and black is so different. I will check it tomorrow.
|
Hi, Ayaka. It would be good if you could give a little bit more detail about this plot:
- What exactly are the "signals"? Are you making a sound somehow? If so, what is producing the sound? What is it's spectrum?
- Are the blue/green/red traces from three different microphones?
- Coherence usually implies a comparison between two signals. Is something being compared in the dashed traces?
- Are the yellow and black traces from different amplifiers?
- What are the units of the Y axis?
|
7591
|
Tue Oct 23 00:41:55 2012 |
Jenne | Update | Alignment | MC locked, spots centered |
[Jenne, Raji]
We replaced the MC Refl path BS with the Y1, as usual, so that the full ~100mW goes to the REFL PD, so we don't have WFS or MC refl camera.
The MC spots were all outside of 1mm, and some were beyond 2mm (for MC1,3, P,Y....MC2 is of course free since we have more DoFs than we need), so we touched (very, very slightly) the zigzag mirrors on the PSL table. We realigned the MC, and now the spots are centered to my satisfaction.
MC1,2,3 Pit, MC1,2,3 Yaw (in mm):
[0.46444020918749457, 8.2634316545130009, -0.41417975237831089, -0.89401481457980592, -0.9323196976382162, -1.543145765853893]
MC2 is way off in pitch, according to this measurement, and it's been consistently going down as we move the MC2 spot in the same direction (up on the monitor), but since we started at +15mm and are now at +8, and we've gone quite a ways, I'm not sure that we really want to go all the way to 0. Anyhow, MC1 and MC3 are the ones which define our input pointing, so we're quitting for tonight.
We will turn on the PZTs and begin with the official vent list for dummies tomorrow. |
Attachment 1: Screenshot-Figure_1.png
|
|
7590
|
Mon Oct 22 21:20:36 2012 |
Jenne | Update | PSL | Beam attenuation optics in place |
[Jenne, Raji (before dinner)]
We put the beam attenuation optics in place. Before putting any optics down, I centered the IOO QPDs, then adjusted the HWPs and PBS such that we remained centered on those QPDs.
Now, I'm about to unblock the beam and let ~100mW into the vacuum so I can lock the MC. Steve and Manasa were putting on the light access connector when I left earlier, so I'm excited to use it! |
7589
|
Mon Oct 22 20:44:49 2012 |
Ayaka | Update | Adaptive Filtering | microphone noise |
I will do some experiments on acoustic noise canceling during my stay.
Now I am planning to cancel acoustic noise from PMC and see how the acoustic noise work and how we should place microphones.
First, I measured the noise in microphones and its circuit.

-blue, green, red, solid lines; microphone signals
-blue, green, red, dashed lines; un-coherent noise in signals
-yellow, black, solid lines; circuit noise (signal input is open, not connected to the microphones)
We can see the acoustic signal above 1 Hz, and the circuit does not seem to limit its sensitivity. But I do not know why yellow and black is so different. I will check it tomorrow. |
7588
|
Mon Oct 22 11:26:12 2012 |
rana | Update | SUS | PITCH damping needed |
Too bad - I thought it would at least give a little damping. Since we want the viscous-like energy loss to be ~49x larger, we need to have the field modulation in the damper (not dumper) increase by ~7. |
7587
|
Mon Oct 22 09:10:07 2012 |
Steve | Update | VAC | @750Torr |
Quote: |
Quote: |
PSL shutter closed, manual block in place, HV turned off. P1 is at 200 Torr now. Jenne is taking over here.
|
Valves closed, 500 torr. Steve will finish off Monday morning, then we'll take off doors and get to work.
|
We are almost at atm. P1 750 Torr, We are slowly reaching equilibrium. |
Attachment 1: ventOct222012.png
|
|
7586
|
Sat Oct 20 20:37:55 2012 |
Den | Update | SUS | PITCH damping needed |
Quote: |
We need stronger magnets if we want eddy current dumping down to Q~1.
|
I've inserted 10mm * 10mm magnets to the 4 corner holes on the front side of the mirror frame according to actuation magnets polarity. I realigned TT and measured Q factor for pitch and yaw, it was 5-10.

I was able to do it for 1 TT only, because others have smaller (~0.1 mm) hole diameter and magnets can't go inside. I tried to warm holes up to 850 F but still was not able to insert a magnet.

|
7585
|
Sat Oct 20 01:23:56 2012 |
Den | Update | SUS | PITCH damping needed |
Quote: |
Measurements confirm that damping using small actuation magnets reduces pitch Q by a factor of 4 and is not enough.
|
I've tested the idea to use coils as eddy current dampers. I terminated them with a wire and measured Q factor during the ringdown test. Sadly, I did not see any significant damping and Q was ~150. We need stronger magnets if we want eddy current dumping down to Q~1.

|
7584
|
Fri Oct 19 19:26:42 2012 |
Jenne | Update | VAC | vent has started |
Quote: |
PSL shutter closed, manual block in place, HV turned off. P1 is at 200 Torr now. Jenne is taking over here.
|
Valves closed, 500 torr. Steve will finish off Monday morning, then we'll take off doors and get to work. |
7583
|
Fri Oct 19 18:30:07 2012 |
Den | Update | SUS | PITCH damping needed |
CQuote: |
From this data I've estimated Q factor for yaw (135 vs 88) and pitch (192 vs 77) (original vs added pitch damping).
|
I've made a more precise measurement of pitch damping using spectrum analyzer.

Measurements confirm that damping using small actuation magnets reduces pitch Q by a factor of 4 and is not enough. |
7582
|
Fri Oct 19 16:45:39 2012 |
Steve | Update | VAC | vent has started |
Quote: |
VENT NOW and FIX ALIGNMENT!
|
PSL shutter closed, manual block in place, HV turned off. P1 is at 200 Torr now. Jenne is taking over here. |
7581
|
Fri Oct 19 16:24:39 2012 |
rana | Update | Locking | Alignment is really bad?? |
VENT NOW and FIX ALIGNMENT! |
7580
|
Fri Oct 19 12:45:12 2012 |
Den | Update | CDS | c1lsc is up after reboot |
|
7579
|
Fri Oct 19 01:21:42 2012 |
Evan | Update | Locking | Aligning PZTs, PRM |
[Evan, Jenne]
Tonight we made an attempt at getting the PRM + ITMY aligned with correct input pointing. We steered the good PZT so that the input beam makes it through the aperture in front of ETMY. We then aligned the PRM so that the retroreflection of the input beam makes it back into the Faraday. After that we tried dithering the alignment of ITMY and the beamsplitter to see if we could see a spot flash across the AS port, but we saw nothing.
For the PRM alignment we set up a camera looking into the window at the Faraday in the IOO chamber; it's called FI_BACK. We stole a 50mm lens from the ETMY face camera.
We also tried looking for beam on IP_POS and IP_ANG. When the input beam is aligned to pass through the ETMY aperture, we can see beam on the steering mirrors preceding IP_POS, but it hits a mirror mount. When the input beam is aligned as it was on Monday, it clips on the ETMY aperture but makes it further along the IP_POS optical path. In both cases, we weren't able to see any beam coming out for IP ANG. |
7578
|
Fri Oct 19 01:11:18 2012 |
Jenne | Update | IOO | MCL, WFS triggers |
Quote: |
Quote: |
I've redone the WFS triggers. I left the MCL trigger alone (for now....I'll come back to it).
The trigger was setup such that (a) it was totally unclear what was going on, by looking at the WFS screen. Koji and I spent some time confused before I remembered that Den did this work recently. Also, for some reason, the triggers were just plain thresholding, not a schmidt trigger, so any time the cavity flashed, the WFS came on. Since the cavity can flash before the mcdown script has a chance to turn off the WFS servos, the outputs of the WFS filters are trying to output thousands of counts, and the signal goes through any time the cavity flashes. Not so good.
|
Your schmitt trigger has 2 threshold values - min and max. Set thresholding value in my trigger to the max of your schmitt trigger and you get the same behavior for MC, triggers are not supposed to turn anything on in this realization as they do for locking with flashing.
|
The problem is that the WFS were being engaged with your triggers every time the MC flashed. That wasn't a schmidt trigger thing, but I like the schmidt trigger better anyway.
Anyhow, it's turned on, and it works really well. It's kind of awesome. I'm really excited to start using the wait block to start pushing even more of the locking out of scripts and into the real time system. |
7577
|
Fri Oct 19 00:55:35 2012 |
Jenne | Update | Computers | c1lsc is down (at least all of the models) |
When Evan and I were dithering the BS and ITMY (see his elog), I noticed that c1lsc was acting weird. the IOP was the only one with the blinky heartbeat. The IOP was all green lights, but all the other models had red for the fb connection, as well as the rightmost indicator (I don't know what that one is for). I logged on to c1lsc and ran 'rtcds restart all'. The script didn't get anywhere beyond saying it was beginning to stop the 1st model (sup, the bottom one on the lsc list). Then all of the cpus went white. I can still ping c1lsc, but I can't ssh to it.
I'm not sure what to do here Jamie. Heelp. |
7576
|
Thu Oct 18 15:36:57 2012 |
Steve | Update | Cameras | Watec cameras & Tamron lenses |
I purchased 3x 1/2" ccd cameras and 3x F 50 mm lenses for the lab.
The spectral sensitivity plot is for an older model 902H. This new model has better sensitivity |
Attachment 1: 10181201.PDF
|
|
7575
|
Thu Oct 18 12:02:32 2012 |
Den | Update | IOO | MCL, WFS triggers |
Quote: |
I've redone the WFS triggers. I left the MCL trigger alone (for now....I'll come back to it).
The trigger was setup such that (a) it was totally unclear what was going on, by looking at the WFS screen. Koji and I spent some time confused before I remembered that Den did this work recently. Also, for some reason, the triggers were just plain thresholding, not a schmidt trigger, so any time the cavity flashed, the WFS came on. Since the cavity can flash before the mcdown script has a chance to turn off the WFS servos, the outputs of the WFS filters are trying to output thousands of counts, and the signal goes through any time the cavity flashes. Not so good.
|
Your schmitt trigger has 2 threshold values - min and max. Set thresholding value in my trigger to the max of your schmitt trigger and you get the same behavior for MC, triggers are not supposed to turn anything on in this realization as they do for locking with flashing. |
7574
|
Thu Oct 18 08:00:40 2012 |
jamie | Update | Computers | Re: Lots of new White :( |
Quote: |
Solved. The power code of c1iscaux was loose.
Has anyone worked around the back side of 1Y3?
I looked into the problem. I went around the channel lists for each slow machines and found the variables are supported by c1iscaux
controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ cd /cvs/cds/caltech/target/c1iscaux
controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ grep C1:IF *
C1IFO_STATE.db:grecord(ai,"C1:IFO-STATE")
It seemed that the machine was not responding to ping. I went to 1Y3 and found the crate was off. Actually this is not correct.
The key was on but the power was off. I looked at the back and found the power code was loose from its inlet.
Once the code was pushed in and the crate was keyed, the white boxes got back online.
Just in case I burtrestored these slow channels by the snapshot at 6:07am on Sunday.
|
I was working around 1Y2 and 1Y3 when I wired the DAC in the c1lsc IO chassis in 1Y3 to the tip-tilt electronics in 1Y2. I had to mess around in the back of 1Y3 to get it connected. I obviously did not intend to touch anything else, but it's certainly possible that I did. |
7573
|
Thu Oct 18 03:57:20 2012 |
Jenne | Update | Locking | Alignment is really bad?? |
The goal of the night was to lock the Y arm. (Since that didn't happen, I moved on to fixing the WFS since they were hurting the MC)
I used the power supplies at 1Y4 to steer PZT2, and watched the face of the black glass baffle at ETMY. (elog 7569 has notes re: camera work earlier) When I am nearly at the end of the PZT range (+140V on the analog power supply, which I think is yaw), I can see the beam spot near the edge of the baffle's aperture. Unfortunately, lower voltages move the spot away from the aperture, so I can't find the spot on the other side of the aperture and center it. Since the max voltage for the PZTs is +150, I don't want to go too much farther. I can't take a capture since the only working CCD I found is the one which won't talk to the Sensoray. We need some more cameras....they're already on Steve's list.
When the spot is a little closer to the center of the aperture than the edge of the aperture (so the full +150V!!), I don't see any beam coming out of AS....no beam out of the chamber at all, not just no beam on the camera. Crapstick. This is not good. I'm not really sure how we (I?) screwed up this thoroughly. Sigh. Whatever ghost REFL beam that Kiwamu and Koji found last week is still coming out of REFL.
Previous PZT voltages, before tonight's steering: +32V on analog power supply, +14.7 on digital. This is the place that the PRMI has been aligned to the past week or so.
Next, just to see what happens, I think I might install a camera looking at the back (output) side of the Faraday so that I can steer PRM until the reflected beam is going back through the Faraday. Team K&K did this with viewers and mirrors, so it'll be more convenient to just have a camera.
Advice welcome. |
7572
|
Thu Oct 18 03:45:56 2012 |
Jenne | Update | IOO | MCL, WFS triggers |
Quote: |
I've added MCL and WFS stop triggers into C1MCS/SUS model. Threshold value of MC_TRANS can be changed in the text entry located in MC2_POSITION medm screen. I tried 2 cases: trigger either blocks signal before MCL filter bank input or after output. Due to filter history in the 1 case MC2 was still slightly disturbed (C1:SUS-MC2_ULPD_VAR ~= 15) right after unlock. In the second case there was no disturbance as we zero output signal, but then I had to add "clear history" command to the mcup script.
WFS triggers block the signal before ASCPIT/YAW filter bank.
|
I've redone the WFS triggers. I left the MCL trigger alone (for now....I'll come back to it).
The trigger was setup such that (a) it was totally unclear what was going on, by looking at the WFS screen. Koji and I spent some time confused before I remembered that Den did this work recently. Also, for some reason, the triggers were just plain thresholding, not a schmidt trigger, so any time the cavity flashed, the WFS came on. Since the cavity can flash before the mcdown script has a chance to turn off the WFS servos, the outputs of the WFS filters are trying to output thousands of counts, and the signal goes through any time the cavity flashes. Not so good.
I have removed the triggering for the angular DoFs from the mcs model (leaving the MCL triggering for now). I have put new triggering into the ioo model, at the error point of the WFS loops. The idea is that if the cavity unlocks, we don't want to lose the current pointing of the mirrors. If the WFS servos were doing a lot of DC work, the bias sliders won't have the full information about where we want the mirrors to point. Since we have the integrators in FM1, removing the input signal should freeze the output signal. I need to modify the WFS on / off script so that this doesn't get turned off every lockloss.
Also, I have implemented (for the first time in a useful model, although I've done some testing in the tst model) the "wait" delay between a cavity locking and the trigger going through. The idea is that we don't necessarily want the WFS to come on simultaneously with the cavity lock. Since the wait delay resets any time it is un-triggered, this also prevents any signals from going through during cavity flashes. The wait block has 3 inputs: (1) a trigger, the output of some kind of trigger block, (2) a number of seconds to wait and (3) the model rate in Hz. The model rate should be set with a constant in the model, the trigger passed from the trigger block, and the wait time in seconds should be available as an epics input.
So far it looks like it's working as I expect, although I'm honestly too tired to do enough testing that I'm satisfied with, so I'm leaving the WFS off for the night. |
7571
|
Thu Oct 18 02:42:38 2012 |
Den | Update | SUS | PITCH damping needed |
Quote: |
I think we can put ø2mm × 10mm long magnetic material inside 4 holes with actuation magnets. Then magnetic field on the other side of the mirror will be close to one produced by actuation magnet. Magnetic cylinder center of inertia will be in the vertical plane where mirror's center of inertia is. So this should not change alignment significantly. Eddy current dumping will be applied to the end of the magnetic cylinder opposite to the magnet using aluminium disks, we have them in the clean room.
|
I've tested this approach. As we do not have required cylinders with high magnetic permittivity, I replaced them with magnets simular to actuator magnets ø2mm × 3mm long. Using them and aluminium disks from other TT I've made a "pitch dumping" construction.

Pitch Q reduced but not that much as I could expect. I did a ringdown test.
Plots:
yaw ringdown using original construction | yaw ringdown with added pitch damping
------------------------------------------------------------------------------------------------------------------------
pitch ringdown using original construction | pitch ringdown with added pitch damping


From this data I've estimated Q factor for yaw (135 vs 88) and pitch (192 vs 77) (original vs added pitch damping). Thess results diverges with the ones obtained by designes. They measured Q~40-50 for original construction. Pitch and yaw have 2 close resonances so this time domain method can not be very precise. I've measured the same with SR785.
 
In these comparison plots excitation was not the same as coils are not plugged in yet, but resonance Q factors can be compared.
|
7570
|
Wed Oct 17 19:35:58 2012 |
Koji | Update | Computers | Re: Lots of new White :( |
Solved. The power code of c1iscaux was loose.
Has anyone worked around the back side of 1Y3?
I looked into the problem. I went around the channel lists for each slow machines and found the variables are supported by c1iscaux
controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ cd /cvs/cds/caltech/target/c1iscaux
controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ grep C1:IF *
C1IFO_STATE.db:grecord(ai,"C1:IFO-STATE")
It seemed that the machine was not responding to ping. I went to 1Y3 and found the crate was off. Actually this is not correct.
The key was on but the power was off. I looked at the back and found the power code was loose from its inlet.
Once the code was pushed in and the crate was keyed, the white boxes got back online.
Just in case I burtrestored these slow channels by the snapshot at 6:07am on Sunday. |
7569
|
Wed Oct 17 18:41:27 2012 |
Jenne | Update | Cameras | Camera looking at ETMY baffle |
The camera titled "watec_mobile" is looking at the front of the black glass baffle (i.e. the side facing the ITM) on the ETMY table. This required (for my quick hacky solution) removing the regular ETMYF camera. Steve has a genius plan (I think) so that we can have both at the same time. Anyhow, eventually we'll move the black glass back, so we'll be back to needing just one camera.
After dinner, I'll try aligning the Yarm. |